# Landauer Principle and Thermodynamics of Computation

### v1

@@ III Landauer Principle considering the open-system dynamics
The study of the dynamics of a quantum system that interacts with other system or systems (environment) is described as an open quantum dynamics (OQD) Breuer et al. (2002); Rivas and Huelga (2012); Rotter and Bird (2015); Viola et al. (1999); Breuer et al. (2016); De Vega and Alonso (2017); Mukhopadhyay et al. (2017); Bhattacharya et al. (2017). It plays a crucial role in almost every aspects of quantum technology as the system which we are interested in is affected by the surrounding noise Breuer et al. (2002).
Till now, we have discussed the LB considering the entropic bounds sans the dynamics during erasing. The dynamics of erasing bear significant importance in practical scenarios. For example, the LB is reached in infinite time. But the practical uses demand finite-time erasure. Also, during the erasing procedure, the environment can get far from its equilibrium (due to finite size), and the fluctuations due to it are crucial for LB. What do the NM features add to the LP? These questions are not only practically important, but can also bring forth fundamental issues. In the following, we review the progress on the finite-time, non-equilibrium, and non-Markovian erasing processes, respectively, mainly in the quantum domain.
@@ III.1 Finite Time Landauer’s Principle
The heat dissipated (LL) in the environment to erase one bit of information is achieved under the assumption of a quasistatic transformation. In practical scenarios, the information erasure occurs in finite time. Therefore, finite-time analysis Andresen (2011) bears much importance both in the classical and quantum regimes. There has been an upsurge of research on finite-time erasure of bits both in the quantum and stochastic thermodynamics Seifert (2012); Van den Broeck and Esposito (2015). In this finite-time regime, the erasing process takes place under non-equilibrium conditions, and the fluctuations in the dissipated heat become significant. This bears important consequences when designing minuscule logical devices that must be able to combat destructive fluctuations that lie well above the LB.
Research on minimizing the average dissipation of a mesoscopic thermodynamic system during finite-time transformations has primarily focused on optimizing a limited (and often small) number of control parameters that influence the system’s potential landscape Schmiedl and Seifert (2007); Bonança and Deffner (2014); Sivak and Crooks (2012); Tafoya et al. (2019); Plata et al. (2020); Bryant and Machta (2020); Boyd et al. (2018); Riechers et al. (2020); Rolandi and Perarnau-Llobet (2023). A significant advancement in this domain was made by Aurell et al. Aurell et al. (2011, 2012), who, using stochastic thermodynamics, developed protocols with full control over the potential landscape to minimize entropy production in both slow and fast limits, while constraining the final state to a fixed microscopic probability distribution. Building on this approach, the authors in their work Proesmans et al. (2020a) propose a framework that also provides full control over the potential landscape but removes the constraint that the final state must be a best-fitted state. Extensions Proesmans et al. (2020b) and alternative approaches Zhen et al. (2021); Dago et al. (2021); Dago and Bellon (2022) in this direction, without accounting for quantum effects, have been explored to establish an optimal bound on the cost of erasure.
@@
Expanding on this, Miller et al. Miller et al. (2020) investigated finite-time erasure under Markovian dynamics generated by the adiabatic Lindblad equation: $\dot{\rho}_{t}=\mathcal{L}\rho_{t}$. The generator $\mathcal{L}$ satisfies quantum details balance condition with respect to the instantaneous control Hamiltonian $H_{t}$ that guarantees an instantaneous fixed point of the dynamics, $\pi_{t}=e^{-\beta H_{t}}/\mbox{Tr}(e^{-\beta H_{t}})$ such that $\mathcal{L}(\pi_{t})=0$. Additionally, the driving of the control Hamiltonian is performed slowly in the time interval $t\in[0,\tau]$, relative to the relaxation timescale of the dynamics which implies that the system remains close to the equilibrium state at all times, i.e., $\rho_{t}=\pi_{t}+\tau^{-1}\delta\rho_{t}$ where corrections of higher orders $O(\tau^{-n})$ can be neglected. In this quasistatic limit, the dissipation approaches the LE bound.
@@
Building on the insights of Miller et al. (2020), recent work Taranto et al. (2023) explored the thermodynamic limits of quantum cooling, juxtaposing LP with Nernst’s unattainability principle. This study introduced a Carnot-Landauer limit, demonstrating that perfect cooling—achieving absolute zero temperature—demands either infinite energy or unbounded control complexity. The cooling bound is given by:
@@
where $\eta_{c}=1-\frac{\beta_{H}}{\beta}$ is the Carnot efficiency, $\beta_{h}$ and $\beta$ being the temperatures of the two thermal baths, $\Delta E_{H}$ quantifies the energy transferred to the subsystem from the hot bath, and $\Delta F_{S}^{(\beta)}$ represents the free energy difference between the initial and final states of the system, and $\epsilon_{E}\geq 0$ denotes the error in the cooling protocol.
@@
The role of coherence in finite-time processes becomes even more pronounced when considering many-body effects and collective erasure. The collective effects of many-body systems in finite-time erasure Rolandi et al. (2023) demonstrated that many-body interactions can significantly reduce dissipated work in finite-time thermodynamic processes. Unlike independent qubit erasure, where excess work scales linearly with system size, many-body protocols exhibit sublinear scaling: $W_{\text{diss}}\propto N^{x},\quad x<1$, indicating an accelerated convergence to LB and an enhancement in the efficiency of information erasure.
@@
Building on Miller’s framework, Van Vu and Saito Van Vu and Saito (2022) further examined the interplay between coherence and erasure speed, establishing that finite-time erasure introduces an additional distance cost (distance error to the ground state) beyond the LL. Their findings reinforced the idea that quantum coherence consistently amplifies dissipation, regardless of the control protocol or driving speed. These insights collectively deepen our understanding of the thermodynamic constraints on quantum information processing, emphasizing the inescapable energy costs imposed by coherence, speed, and many-body effects.
@@
Importantly, one can implement an arbitrarily fast erasure process through a highly optimized algorithm that prescribes the exact microstate manipulations needed to transform any initial state into a desired pure final state. If algorithmic complexity is unconstrained, meaning we are allowed arbitrarily complex and precise instructions (programs), then it is in principle possible to design a perfect erasure protocol that acts optimally on any input, driving the system to a unique pure state in finite time. These algorithms would encode a complete understanding of the initial conditions and the exact transformations required, bypassing the usual thermodynamic cost that arises from ignorance or randomness in the input. However, this does not violate Landauer’s principle; rather, it circumvents the physical cost by assuming unbounded computational resources, shifting the “cost” from thermodynamic to computational complexity  Zurek (1989a).
@@
In practice, limitations arise due to thermodynamic speed limits Deffner and Campbell (2017) and coherence-induced excess dissipation Faist et al. (2015), finite computational resources. Recent studies Van Vu and Saito (2022) further quantify the finite-time cost, revealing that even with optimal control, a residual energy cost remains due to quantum fluctuations and irreversibility. Thus, while the Landauer limit can be theoretically approached, the interplay of coherence, control complexity, and finite-time constraints fundamentally prevents its exact realization in realistic settings.
III.2 Non-equilibrium Landauer Process
@@
The miniaturization in modern technology has led to the development of small systems that are out-of-equilibrium in classical Jarzynski (2011); Seifert (2012) as well as in the quantum regime Esposito et al. (2009); Campisi et al. (2011a); Goold et al. (2016). The fluctuation relation Jarzynski (2011, 1997c, 2004); Crooks (1999); Tasaki (2000); Kurchan (2000); Mukamel (2003); Campisi et al. (2011b) plays a promising role in understanding the thermodynamics of these small systems that operate under the non-equilibrium condition where the thermal and quantum fluctuations cannot be neglected. In this review article, we specifically focus on nonequilibrium erasure protocols within the quantum regime.
A key challenge in non-equilibrium thermodynamics is characterizing the dissipation associated with quantum erasure. Reeb and Wolf Reeb and Wolf (2014) established a fundamental lower bound for heat dissipation in equilibrium conditions, highlighting the role of quantum coherence in modifying LP. However, real-world erasure processes often operate in non-equilibrium regimes, where such equilibrium-based bounds may not be directly applicable. To bridge this gap, their framework has been recast from a fluctuation relation perspective, which provides a generalized thermodynamic bound for erasure beyond equilibrium.
In Goold et al. (2015), an erasure protocol that involves a finite-size environment that interacts with the system is proposed. The system is a single qubit system that is coupled to a finite-dimensional environment. A tantalizing fact is that the nonunitality of the open system dynamics leads to a tighter bound for the heat dissipation for the erasure process. This opened the door for the analysis of the cost of computation in non-equilibrium circumstances. Following the same methodology, a comparative analysis of the performance of LB with that of the bound proposed in Goold et al. (2015) for the non-equilibrium condition is addressed Campbell et al. (2017); Zhang et al. (2023). Reeb and Wolf Reeb and Wolf (2014) in their work provided the minimum framework that is required for the Landauer process in the equilibrium condition. The minimum framework for the execution of the Landauer process in the non-equilibrium condition has been addressed in Taranto et al. (2018), where the system and the environment are equivalent to the models in Campbell et al. (2017); Zhang et al. (2023). Although this approach offers valuable insights, it remains limited to particular models and does not yet provide a universally applicable methodology. To rigorously understand LP in the context of quantum non-equilibrium dynamics, further investigation is essential to develop a minimal yet broadly generalizable framework that captures the complexities unique to quantum systems.
The prerequisites for the erasure protocol are (i) A system $\mathcal{S}$ with $H_{\mathcal{S}}$ as the free Hamiltonian of the system. (ii) An environment $\mathbf{E}$ initially uncorrelated with the system, i.e., $\rho_{\mathcal{S}\mathbf{E}}=\rho_{\mathcal{S}}\otimes\rho_{\mathbf{E}}$. (iii) The initial state of the environment will be the Gibbs state. (iv) The interaction of the environment and the system is unitary. The heat fluctuation relation of the environment for the erasure protocol is:
@@
which follows from the equality proposed by Jarzynski Jarzynski (1997c) using the work distribution. $P(Q)$ denotes the heat distribution of the environment. ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}_{l=jk}}=\sqrt{\lambda_{j}}\langle s_{k}|\hat{U}|s_{j}\rangle$, with $\lambda_{j}$ being the eigenvalue and $|s_{j}\rangle$ being the eigenstate of $\rho_{\mathcal{S}}$. The process is unital iff $\hat{\textbf{A}}=\mathbf{I}_{\mathbf{E}}$. If the operator $\hat{\textbf{A}}$ is expanded in terms of the initial state of both $\mathcal{S}$ and $\mathbf{E}$ under the action of the global unitary evolution, the heat fluctuation relation becomes
$\langle e^{-\beta Q}\rangle=\text{tr}[\hat{\textbf{M}}\rho_{\mathcal{S}}],$ (7)
where $\hat{\textbf{M}}=\text{tr}_{\mathbf{E}}[\hat{U}^{\dagger}\mathbf{I}_{\mathcal{S}}\otimes\rho_{\mathbf{E}}\hat{U}$]. The desired heat dissipation during the erasure process is
@@
where $\mathbf{B}_{Q}=-\,\ln(\text{tr}[\hat{\textbf{A}}\rho_{\mathbf{E}}])=-\,\ln(\text{tr}[\hat{\textbf{M}}\rho_{\mathcal{S}}])$ is termed the thermodynamic bound. The operator $\hat{\textbf{A}}$ is dependent on the choice of the state of the system. So, for the execution of the erasure process of a state of choice, one needs to compute $\hat{\textbf{A}}$ for every instance. On the other hand, $\hat{\textbf{M}}$ as defined in Eq. (7) can be easily evaluated just by implementing the unitary interaction. This proposed bound is tested on a physical system where the system is considered to be a single qubit and the environment comprises an interacting spin chain (as shown in Fig. 1). The Hamiltonian of the environment, given by the $XX$ model, is
@@
Here $\phi=\cos(2Jt)$ at time $t$, $p=[1+\tanh(\beta B)]/2$ and $\alpha\in\mathbb{R}$. For $\alpha\approx 1$ and for a particular temperature it is observed that $\mathbf{B}_{Q}\geq\mathbf{B}_{rw}$, where $\mathbf{B}_{rw}$ denotes the bound proposed in Reeb and Wolf (2014).
@@
A comparison of different forms of the LBs reveals a subtle dependence on the initial state of the system and the environment’s temperature. In the parameter space, sharp boundaries emerged when assessing the relative effectiveness of these bounds Campbell et al. (2017). There were scenarios where the bounds were negative in both frameworks, yet the process’s heat dissipation remained positive. In such cases, the bounds proved to be weaker than the Clausius statement of the second law.
The full counting statistic Esposito et al. (2009) formalism, when applied to analyze the bounds (lower and upper) on the average heat dissipation in an erasure process, a single-parameter bound is developed that can be made arbitrarily tight and is independent of the map being used for the execution of the process Guarnieri et al. (2017). This inherently marks the difference of the lower bound of this process over the previous bound Goold et al. (2015). For this formalism, the minimal set of assumptions is considered as follows in Reeb and Wolf (2014), which validates LP.
The full counting statistics of heat dissipation are defined as the change of energy in the environment that characterizes the mean value of heat dissipation. The lower bound for the heat dissipation using the CGF Rockafellar (1970) is
@@
with $\eta$ being the counting parameter. Using the large deviation function (useful for studying dynamical phase transitions), which is a powerful tool to study statistical properties for a long time scale Garrahan and Lesanovsky (2010); Lesanovsky et al. (2013); Pigeon et al. (2015), the upper bound for heat dissipation is proposed as
@@
The Landauer-like bound here is proposed with one parameter on a two-time measurement protocol. For $\eta=\beta$ the derived form of the bound is equivalent to that of the bound proposed in Goold et al. (2015). This bound is tested on a physical system where the system is considered to be a three-level V-system and the environment is modeled by a two-level system (as in Fig. 2). The transition is pumped by a transition frequency $\Omega_{1}$.
@@
This physical model highlights the tightness of the proposed bound for heat dissipation during the execution of the erasure protocol.
III.3 Landauer Bound in Non-Markovian Process
The Markovian approximation is the most convenient way to express the dynamics of OQS. In this process, the evolution timescale is considered to be larger than the correlation time of the environment. In other words, in this approximation, the memory effect (or information backflow) is neglected. However, it plays a non-minuscule role in the analysis of the dynamics of the system and thermodynamic cost. There comes the role of non-Markovian (NM) dynamics to explore the dynamics of the system De Vega and Alonso (2017); Breuer et al. (2016). Various methods and approaches have surfaced for defining as well as quantifying non-Makovianity Breuer et al. (2009); Rivas et al. (2010); Chruściński and Maniscalco (2014).
Out of the various methods, one is the applicability of the collision model for NM dynamics Rau (1963); Alicki and Lendi (2007); Scarani et al. (2002); Ziman and Bužek (2005); Gennaro et al. (2009). In the collision model, the NM dynamics for a system can be achieved when the interaction of the system and the environment is intervened by degrees of freedom of the ancilla. Under NM dynamics, where coherence endures and information retrogrades from the environment, a coherence-dependent correction to LB Reeb and Wolf (2014) becomes imperative. Unlike Markovian erasure, where dissipation follows a strict lower bound, NM effects can modify these limits, potentially reducing or amplifying heat dissipation. Here, we extend Landauer’s principle to NM dynamics, exploring how memory effects influence erasure costs and coherence-driven corrections. This provides a more comprehensive view of quantum erasure beyond the standard Markovian framework.
The formulation of a Ladauer-like principle for heat flux in an erasure process delineated by the collision model of OQS is first proposed in Lorenzo et al. (2015). The process of analyzing the information to energy conversion in the collision model provides the platform for the foundation of the LP in NM dynamics.

### v2

@@ III Landauer Principle considering the open-system dynamics
The study of the dynamics of a quantum system that interacts with other system(s) (environment) is described as an open quantum dynamics Breuer et al. (2002); Rivas and Huelga (2012); Rotter and Bird (2015); Viola et al. (1999); Breuer et al. (2016); De Vega and Alonso (2017); Mukhopadhyay et al. (2017); Bhattacharya et al. (2017). It plays a crucial role in almost every aspects of quantum technology as the system which we are interested in is affected by the surrounding noise Breuer et al. (2002).
Till now, we have discussed the LB considering the entropic bounds sans the dynamic detail or fluctuations during the erasing processes. The dynamics of erasing bear significant importance in practical scenarios. For example, LB is reached in infinite time. But the practical uses demand finite-time erasure. Also, during the erasing procedure, the environment can get far from equilibrium (due to finite size), and the fluctuations due to it are crucial for LB. What do the NM features add to the LP? These questions are not only practically important, but can also bring forth fundamental issues. In the following, we review the progress on the finite-time, non-equilibrium, and non-Markovian erasing processes, respectively, mainly in the quantum domain.
@@ III.1 Finite Time Landauer’s Principle
The heat dissipated in the environment to erase one bit of information is achieved under the assumption of a quasistatic transformation. In practical scenarios, the information erasure occurs in a finite time. Therefore, the finite-time analysis Andresen (2011) bears much importance both in the classical and quantum regimes. There has been an upsurge of research on finite-time erasure of bits both in the quantum and stochastic thermodynamics Seifert (2012); Van den Broeck and Esposito (2015). In this finite-time regime, the erasing process takes place under non-equilibrium conditions, and the fluctuations in the dissipated heat become significant. This bears important consequences when designing minuscule logical devices that must be able to combat destructive fluctuations that lie well above the LB.
Research on minimizing the average dissipation of a mesoscopic thermodynamic system during finite-time transformations has primarily focused on optimizing a limited (and often small) number of control parameters that influence the system’s potential landscape Schmiedl and Seifert (2007); Bonança and Deffner (2014); Sivak and Crooks (2012); Tafoya et al. (2019); Plata et al. (2020); Bryant and Machta (2020); Boyd et al. (2018); Riechers et al. (2020); Rolandi and Perarnau-Llobet (2023). A significant advancement in this domain was made by Aurell et al. Aurell et al. (2011, 2012), who, using stochastic thermodynamics, developed protocols with full control over the potential landscape to minimize entropy production in both slow and fast limits, while constraining the final state to a fixed microscopic probability distribution. Building on this approach, the authors in their work Proesmans et al. (2020a) proposed a framework that enables full control over the system’s potential without requiring a pre-specified final state, allowing more flexible and efficient erasure protocols. Using this, they derive tight lower and upper bounds on the dissipated work during optimal finite-time bit erasure. Extensions Proesmans et al. (2020b) in the direction of Proesmans et al. (2020a) and alternative approaches Zhen et al. (2021); Dago et al. (2021); Dago and Bellon (2022) in this direction, without accounting for quantum effects, have been explored to establish an optimal bound on the cost of erasure.
@@
Expanding on this, Miller et al. Miller et al. (2020) investigated finite-time erasure under Markovian dynamics generated by the adiabatic Lindblad equation: $\dot{\rho}_{t}=\mathcal{L}\rho_{t}$. The generator $\mathcal{L}$ satisfies quantum details balance condition with respect to the instantaneous control Hamiltonian $H_{t}$ that guarantees an instantaneous fixed point of the dynamics, $\pi_{t}=e^{-\beta H_{t}}/\mbox{Tr}(e^{-\beta H_{t}})$ such that $\mathcal{L}(\pi_{t})=0$. Additionally, the driving of the control Hamiltonian is performed slowly in the time interval $t\in[0,\tau]$, relative to the relaxation timescale of the dynamics which implies that the system remains close to the equilibrium state at all times, i.e., $\rho_{t}=\pi_{t}+\tau^{-1}\delta\rho_{t}$ where corrections of higher orders $O(\tau^{-n})$ can be neglected. In this quasistatic limit, the dissipation approaches the LB.
@@
It is important to mention that perfectly cooling a system to its desired final state and perfect erasing of information are fundamentally equivalent Taranto et al. (2023), as both correspond to reducing the rank of the system to 111. However, Nearst’s unattainability principle, often referred to as the third law of thermodynamics, states that to perfectly cool a physical system, one needs diverging resources, either in terms of energy or time. The authors of Taranto et al. (2023) showed that it is possible to cool perfectly with finite energy and time provided one is equipped with coherent control whose complexity diverges. In essence, under coherent control, perfect cooling is possible if either time, energy, or control complexity diverges. They further showed that with incoherent control (or cooling aided by a heat engine), perfect cooling remains fundamentally unattainable within finite time and finite control complexity, regardless of the amount of energy extracted from the auxiliary engine.
Consequently, they established that for cooling under incoherent control with finite control complexity, it is possible to cool a system arbitrarily close to a desired temperature within accuracy $\epsilon_{\beta}>0$, with infinite number of operations and thereby in infinite time, with a energy, as measured by the amount of heat drawn from a hot bath at inverse temperature $\beta_{h}$, which is arbitrarily close to the bound called Carnot-Landauer limit, that is
@@
where $\eta_{c}=1-\frac{\beta_{H}}{\beta}$ is the Carnot efficiency, $\Delta E_{H}$ quantifies the energy drawn from the hot bath, and $\Delta F_{S}^{(\beta)}$ represents the free energy difference between the initial and final states of the system at inverse temperature $\beta$, and $\epsilon_{E}>0$. The Carnot-Landauer bound $\frac{1}{\eta_{c}}\Delta F_{S}^{(\beta)}$ reduces to the LB when $\beta_{H}=0$. Both these bounds can be reached asymptotically with diverging control complexity or time.
@@
The role of quantum features like coherence in finite-time processes becomes even more pronounced when considering many-body effects and collective erasure. The collective effects of many-body systems in finite-time erasure Rolandi et al. (2023) demonstrated that many-body interactions can significantly reduce dissipated work in finite-time thermodynamic processes. Unlike independent qubit erasure, where excess work scales linearly with system size, many-body protocols exhibit sublinear scaling: $W_{\text{diss}}\propto N^{x},\quad x<1$, indicating an accelerated convergence to LB and an enhancement in the efficiency of information erasure.
@@
In this direction, Van Vu and Saito Van Vu and Saito (2022) further examined the interplay between coherence and erasure speed, establishing that finite-time erasure introduces additional cost dependent on the distance error to the desired final state and inversely proportional to the total time taken by the process, beyond the standard LB. Their findings reinforced the idea that quantum coherence consistently amplifies dissipation, regardless of the control protocol or driving speed. These insights collectively deepen our understanding of the thermodynamic constraints on quantum information processing, emphasizing the inescapable energy costs imposed by coherence, speed, and many-body effects.
@@
Importantly, one can implement an arbitrarily fast erasure process through a highly optimized algorithm that prescribes the exact microstate manipulations needed to transform any initial state into a desired pure final state. If algorithmic complexity is unconstrained, meaning we are allowed arbitrarily complex and precise instructions (programs), then it is in principle possible to design a perfect erasure protocol that acts optimally on any input, driving the system to a unique pure state in finite time. These algorithms would encode a complete understanding of the initial conditions and the exact transformations required, bypassing the usual thermodynamic cost that arises from ignorance or randomness in the input. However, this does not violate LP; rather, it circumvents the physical cost by assuming unbounded computational resources, shifting the “cost” from thermodynamic to computational complexity  Zurek (1989a).
@@
In practice, limitations arise due to thermodynamic speed limits Deffner and Campbell (2017) and coherence-induced excess dissipation Faist et al. (2015), finite computational resources. While the LB can be theoretically approached, the interplay of coherence, control complexity, and finite-time constraints fundamentally prevents its exact realization in realistic settings.
III.2 Non-equilibrium Landauer Erasure
@@
The miniaturization in modern technology has led to the control of small systems that are out-of-equilibrium in classical Jarzynski (2011); Seifert (2012) as well as in the quantum regime Esposito et al. (2009); Campisi et al. (2011a); Goold et al. (2016). The fluctuation relation Jarzynski (2011, 1997c, 2004); Crooks (1999); Tasaki (2000); Kurchan (2000); Mukamel (2003); Campisi et al. (2011b) plays the central role in understanding the thermodynamics of these small systems that are far from the equilibrium. Here, we mainly focus on nonequilibrium erasure protocols within the quantum regime.
A key challenge in non-equilibrium thermodynamics is characterizing the dissipation associated with quantum erasure. Reeb and Wolf Reeb and Wolf (2014) have established a fundamental lower bound for heat dissipation, using the tools of quantum information theory and quantum statistical mechanics, that includes corrections arising from finite-size reservoir and system-bath correlation. Though Reeb and Wolf have not incorporated the fluctuations arising due to the finite-size effects in their derivation, one might be tempted to think that the correction due to the finite-size effects is somewhat related to the nonequilibrium fluctuations, as we are far away from the thermodynamic limit. Indeed, an equivalent form Eq. (2), has been derived independently in Esposito et al. (2010) to express irreversible entropy production of a finite system out of equilibrium. This indicates the necessity of LB-like bound in far from equilibrium conditions, considering the fluctuations explicitly in the setup. To bridge this gap, researchers have sought to establish generalized bounds for information erasure beyond equilibrium, employing the concepts of fluctuation relations Jarzynski (2011, 1997c, 2004).
In Goold et al. (2015), an erasure protocol that involves a finite-size environment is studied, and a modified bound considering the fluctuations during the protocol is proposed. This has opened the door for the analysis of the cost of computation in non-equilibrium circumstances. Under the same four assumptions that have been used in Reeb and Wolf (2014) to derive Eq. (2), the authors of Goold et al. (2015) have first derived the following heat fluctuation-like relation of the environment for the erasure protocol
@@
Here $P(Q)$ denotes the heat distribution of the environment, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}A_{l=jk}}=\sqrt{\lambda_{j}}\langle s_{k}|U|s_{j}\rangle$, with $\lambda_{j}$ being the eigenvalue and $|s_{j}\rangle$ being the eigenstate of $\rho_{\mathcal{S}}$, $U$ is the joint unitary operator acting on the system and environment and $\textbf{A}=\sum_{l}A_{l}A^{\dagger}_{l}$. The above equation is in the same spirit of Jarzynski equality which connects the average exponentiated work to the equilibrium free energy exploiting the work distribution Jarzynski (1997c). If the operator A is expanded in terms of the initial state of both $\mathcal{S}$ and $\mathbf{E}$ under the action of the global unitary evolution, the heat fluctuation relation becomes
$\langle e^{-\beta Q}\rangle=\text{tr}[\textbf{M}\rho_{\mathcal{S}}],$ (7)
where $\textbf{M}=\text{tr}_{\mathbf{E}}[U^{\dagger}\mathbf{I}_{\mathcal{S}}\otimes\rho_{\mathbf{E}}U$]. The desired heat dissipation during the erasure process satisfies111 In this nonequilibrium LP section, we denote the averaged heat inside bra-ket to distinguish it from the heat exchange in a single trajectory. In Eq. (1), Eq. (2), and all other sections of the review, $Q$ (“heat”) denotes heat dissipated on average.
@@
where $\mathbf{B}_{Q}=-\,\ln(\text{tr}[\textbf{A}\rho_{\mathbf{E}}])=-\,\ln(\text{tr}[\hat{\textbf{M}}\rho_{\mathcal{S}}])$, termed as the thermodynamic bound in Campbell et al. (2017). The process is unital iff $\textbf{A}=\mathbf{I}_{\mathbf{E}}$, and therefore the bound can be insightful to compare the heat fluctuations of the environment under unital and nonunital channels. The operator A is dependent on the choice of the state of the system. So, for the execution of the erasure process of a state of choice, one needs to compute A for every instance. On the other hand, M as defined in Eq. (7) can be evaluated just by implementing the unitary interaction.
This proposed bound is tested Goold et al. (2015) on a physical system where the system is considered to be a single qubit and the environment comprises an interacting spin chain (as shown in Fig. 1). The Hamiltonian of the environment, given by the $XX$ model, is
@@
Here $\phi=\cos(2Jt)$ at time $t$, $p=[1+\tanh(\beta B)]/2$ and $\alpha(\in\mathbb{R})$ is an initial state parameter of the system. For $\alpha\approx 1$ and for a particular temperature it is observed that $\mathbf{B}_{Q}\geq\mathbf{B}_{rw}$, where $\mathbf{B}_{rw}$ denotes the bound proposed in Reeb and Wolf (2014). It is important to mention that the tightness of the bound $\mathbf{B}_{Q}$ over $\mathbf{B}_{rw}$ is parameter dependent. There can be regions in the parameter space where the bound $\mathbf{B}_{rw}$ is tighter than the thermodynamic one.
@@
In Campbell et al. (2017), the authors have performed a detailed comparative study of the thermodynamic bound $\textbf{B}_{Q}$ and the bound $\beta\langle Q\rangle\geq\Delta S$ (see Eq. (2)), dubbed as the entropic bound. They have explored the role of the initial state parameters on the tightness of the two bounds with respect to the dissipated heat. It has also been shown that the thermodynamic bound is not sensitive to the initial state coherence, in contrast to the entropic bound. Moreover, they have shown that there are regions where both bounds are ineffective, being less informative than the Clausius statement of the second law.
In Taranto et al. (2018), the authors have reported a very important aspect of heat fluctuations in erasure protocol. Building upon the framework of Goold et al. (2015), they have shown using generalized concentration of measure that fluctuations of heat from its mean are exponentially damped with the size of the larger subsystem (system or environment), and this suppression scales linearly with increasing inverse temperature of the environment, regardless of the system’s microscopic dynamics.
Though Goold et al. (2015) has paved the way to incorporate the fluctuations-like relations of heat to get an improved bound on the dissipated heat, there remains plenty of scope for further explorations. For example, the quest for a bound that considers fluctuations of heat, which is more informative than the Clausius inequality in general, remains. Generalization of these bounds to an initially correlated system-environment would be challenging and intriguing.
The full counting statistic Esposito et al. (2009) formalism has been applied to analyze the bounds (lower and upper) on the average heat dissipation in an erasure process in Guarnieri et al. (2017). A single-parameter family of bounds is developed that can be made arbitrarily tight and is independent of the underlying map used for the execution of the process Guarnieri et al. (2017). This inherently marks the difference of the lower bound of the nonequilibrium erasure process over the previous bound Goold et al. (2015). For this formalism, the minimal set of assumptions is considered as in Reeb and Wolf (2014).
In this full counting statistics of dissipated heat formalism, the lower bounds for the heat dissipation can be given using the (CGF) Rockafellar (1970) as
@@
$\eta$ being the counting parameter. The Landauer-like bounds here are proposed with one parameter on a two-time measurement protocol. For $\eta=\beta$ the result of Goold et al. (2015) is recovered.
Furthermore, using the large deviation function, which is a powerful tool to study statistical properties for a long time scale Garrahan and Lesanovsky (2010); Lesanovsky et al. (2013); Pigeon et al. (2015), the upper bounds of heat dissipation are given as
@@
These bounds are tested on a physical system where the system is considered to be a three-level V-system and the environment is modeled by a two-level system (as in Fig. 2). The transition is pumped by a transition frequency $\Omega_{1}$.
@@
This physical model highlights the tightness of the proposed bounds.
In Zhang et al. (2023), the authors have studied thermodynamics of nonequilibrium reservoirs with the help of Landauer-like bound. They have shown that a bound in terms of equality, called the Landauer-like bound, always holds even for nonequilibrium reservoirs, but an inequality obtained from the equality by excluding a few contributions might be violated. Violation of this inequality, dubbed as Landauer-like inequality, indicates the possibility of spontaneous heat transfer from a cold bath to a hot bath, beating the Carnot bound and extracting work from a single nonequilibrium reservoir.
III.3 Landauer Bound under Non-Markovian Dynamics
The Markovian approximation is the most convenient way to express the dynamics of OQS Rau (1963); Alicki and Lendi (2007); Scarani et al. (2002); Ziman and Bužek (2005); Gennaro et al. (2009). In this process, the evolution timescale of the system is considered to be larger than the correlation time of the environment. In other words, in this approximation, the memory effect (or information backflow) is neglected. However, it plays a non-minuscule role in the analysis of the dynamics of the system and thermodynamic cost of erasure. This backflow of information or non-Markovian (NM) De Vega and Alonso (2017); Breuer et al. (2016) dynamics has been studied following various methods and approaches Breuer et al. (2009); Rivas et al. (2010); Chruściński and Maniscalco (2014).
Under NM dynamics, where coherence endures and information retrogrades from the environment, a coherence-dependent correction to LB Reeb and Wolf (2014) becomes imperative. Unlike Markovian erasure, where dissipation follows a strict lower bound, NM effects can modify these limits, potentially reducing or amplifying heat dissipation. Here, we discuss LP under NM dynamics, exploring how memory effects influence erasure costs and coherence-driven corrections. This provides a more comprehensive view of quantum erasure beyond the standard Markovian framework.
Out of the various methods for exploring NM dynamics, one of the prominent approaches is adoption of the collision model for OQS Rau (1963); Alicki and Lendi (2007); Scarani et al. (2002); Ziman and Bužek (2005); Gennaro et al. (2009). The process of analyzing the information to energy conversion in the collision model provides the platform for the foundation of the LP under NM dynamics. The formulation of a Ladauer-like principle for heat flux in an erasure process delineated by the collision model has been first proposed in Lorenzo et al. (2015).